# Know What You Don’t Know: Uncertainty Calibration of Process Reward Models

### v1

@@ 3.1 Instance-Adaptive Inference-Time Scaling
Existing inference-time scaling methods typically allocate a fixed sampling budget $N$—for example, by generating $N$ complete reasoning trajectories in BoN. However, this approach can be inefficient: for challenging tasks, $N$ samples may be insufficient to produce a correct answer, while for easier ones, it may waste computation. Just as humans allocate more effort to harder problems and less to simpler ones, it is desirable for LLMs to adjust their compute usage adaptively. Note that in the present work, we (perhaps erring on the side of optimism) assign the maximum budget to the hardest problems, but in principle we could also choose to say “I don’t know” or route to a more capable model if the probability of success is too low (difficulty too high).
@@
Fix a question–prefix pair and let $p\in[0,1]$ be the probability, conditioned on the current generation, that a single continued trajectory sampled from the language model is correct. For a target probability $C\in(0,1)$ we define the independent-sampling sample-complexity
$N^{\star}(p,C)\;\triangleq\;\min\Bigl{\{}n\in\mathbb{N}:\;\Pr\bigl{(}\text{at least one out of $n$ trajectories is correct}\bigr{)}\geq C\Bigr{\}}.$
@@
However, as demonstrated in Section 5.1, PRMs frequently overestimate LLMs’ success probabilities—particularly for weaker models or on hard, out-of-distribution queries (Figure 2). Relying on these inflated scores leads to suboptimal inference-time scaling by selecting an excessively small $N_{\mathrm{IAS}}$, whereas calibration mitigates these pitfalls and enables principled, cost-effective decision making.
4 PRM Calibration and Instance-Adaptive Scaling
Now, we present a calibration-via-finetuning approach to mitigate the aforementioned issues, along with our proposed instance-adaptive inference-time scaling methods based on calibrated rewards.
4.1 Uncertainty-Aware Calibration through Quantile Regression
To evaluate the calibration quality of the given LLM and PRM, we construct a validation set [^1] Our calibration datasets—consisting of triplets of questions, prefixes, and success-probability estimates—are available at https://huggingface.co/datasets/young-j-park/prm_calibration. $\cup_{q\in\mathcal{Q}}\{\mathbf{x}^{(q,1)},\mathbf{x}^{(q,2)},\dots,\mathbf{x}^{(q,N_{\mathrm{val}})}\}$ where $\mathcal{Q}$ is a validation question set. For simplicity, we omit the superscript $q$ in the following discussion. First, for each question $q\in\mathcal{Q}$, we generate $N_{\mathrm{val}}$ independent reasoning trajectories using the LLM: $\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},\dots,\mathbf{x}^{(N_{\mathrm{val}})}\}.$ Next, for each reasoning trajectory $\mathbf{x}^{(i)}$, and each of its prefix trajectories $x_{1:t}^{(i)}$, we generate $N_{\mathrm{MC}}$ additional follow-up trajectories conditioned on both the question and the prefix. By evaluating and counting the number of correct follow-up trajectories, $Z^{(i,t)}$, we estimate the reward and the success probability as: $\hat{r}^{(i,t)}=\mathsf{PRM}_{\phi}(q,x_{1:t}^{(i)})$ and $\tilde{p}^{(i,t)}\approx Z^{(i,t)}/N_{\mathrm{MC}}$, respectively. Please see details in Appendix C.1.
We seek to fine-tune PRMs to predict these success rates $p^{(i,t)}$ given the input and generated steps so far $x_{1:t}^{(i)}$. Since calibration data is non-exhaustive and model capacity is finite, as with any machine learning model, there must always be some uncertainty in the prediction of $p^{(i,t)}$. Because of ERM training, in essence, these models seek to make their best loss-minimizing guess, typically something akin to the conditional mean. [^2] Exactly the conditional mean for MSE loss and approximately so for CE loss if the uncertainty is narrow. This behavior is insufficient for our needs—if we are to use the model for instance-adaptive scaling, the required number of samples we compute will be too small for a high fraction of the test cases (those whose true $p$ are smaller than the conditional mean). For IAS, we would instead like to have a lower bound on the true $p$, for at least a large fraction of the realizations. In other words, we would like to predict a (left) quantile of the posterior distribution for $p$.
To achieve this, we modify the loss to perform quantile regression. Rather than fine-tuning PRMs solely with deterministic error, we propose a modification to the reward model. Specifically, we adjust the output dimension of the prediction head to generate multiple predictions corresponding to different quantiles (e.g., 10%, 50%, and 90%) rather than a single reward value. Details of this modification and the initialization process for the altered prediction head are described in Appendix C.3. Finally, we fine-tune the modified model using weighted quantile loss (wQL):
$\mathsf{wQL}(\hat{r},\tilde{p})\triangleq\frac{1}{N_{q}}\sum_{n=1}^{N_{q}}\left[\beta_{n}\cdot\max\left(0,\;\tilde{p}-\hat{r}^{(\beta_{n})}\right)+(1-\beta_{n})\cdot\max\left(0,\;\hat{r}^{(\beta_{n})}-\tilde{p}\right)\right],$
where $N_{q}$ denotes the number of quantiles, and $\hat{r}^{(\beta_{n})}$ represents the $\beta_{n}$-quantile of the prediction $\hat{r}$.
@@ 4.2 IAS: Calibrated Reward-based Instance-Adaptive Scaling
BoN+IAS.
Rather than drawing a fixed sample size $N$ for best-of-$N$ decoding, our IAS framework adaptively decides the number of trajectories to generate:
$N_{\mathrm{IAS}}\triangleq\min\{n\mid N_{\mathrm{IAS}}(\hat{r}^{(\beta)},C)\leq n,\quad 1\leq n\leq N\}=\min\{\lceil N_{\mathrm{IAS}}(\hat{r}^{(\beta)},C)\rceil\;,\;N\}.$
BS+IAS-of-$M$ (with fixed $K$).
Suppose at a given beam-search step we have $K$ candidate prefixes, each with a minimum success probability (or calibrated lower bound) $\hat{r}^{(\beta)}_{\mathrm{min}}$. To ensure at least one correct completion among these $K$ prefixes, IAS expands the beam per prefix by:
$M_{\mathrm{IAS}}\triangleq\min\{m\mid N_{\mathrm{IAS}}(\hat{r}_{\mathrm{min}}^{(\beta)},C)\leq K\times m,\quad 1\leq m\leq M\}=\min\Bigl{\{}\Bigl{\lceil}\tfrac{N_{\mathrm{IAS}}\!\bigl{(}\hat{r}^{(\beta)}_{\mathrm{min}},C\bigr{)}}{\,K\,}\Bigr{\rceil}\;,\;M\Bigr{\}}.$
BS+IAS-of-$K$ (with fixed $M$).
In beam search with a fixed number of expansions per leaf $M$, if we want to preserve only $K_{\mathrm{IAS}}$ among $N$ expanded prefixes, IAS adaptively prunes to
$K_{\mathrm{IAS}}\triangleq\min\{k\mid N_{\mathrm{IAS}}(\hat{r}_{k}^{(\beta)},C)\leq k\times M,\quad 1\leq k\leq K\},$
where $\hat{r}_{k}^{(\beta)}$ denotes the $k$-th highest calibrated lower bound among the $N$ candidate prefixes. This ensures that the resulting algorithm can collectively draw $K_{\mathrm{IAS}}\times M\;\geq\;N_{\mathrm{IAS}}\!\bigl{(}\hat{r}_{K_{\mathrm{IAS}}}^{(\beta)},C\bigr{)}$ samples, keeping the chance of getting at least one correct completion above $C$.
@@
Set $N=\infty$. Let $\hat{r}^{(\beta)}$ be the prediction of the $\beta$th quantile, and suppose we have held out an $n$-sample validation set $\mathcal{V}_{n}$. Then, on test input $X_{n+1}$
$P(\text{success best-of-}N_{\mathrm{IAS}}(\hat{r}^{(\beta)},C)|X_{n+1})\geq C(1-\tilde{\beta}),$
for some $\tilde{\beta}\geq\beta$ depending on $\mathcal{V}_{n}$ (see Appendix B.2.1).
@@ 5.1 Fine-Tuning PRMs for Better Calibration
We evaluate calibration errors of off-the-shelf PRMs and then show how our fine-tuning strategy reduces these errors. To start with, we construct a calibration dataset by randomly sampling 500 questions from the MATH training set. We assess calibration performance using standard metrics: Brier score, positive Brier score (i.e., the mean square of overestimation error, $\max\{\hat{y}-y,0.0\}$), ECE, and ACE between ground-truth Monte Carlo success rates and the PRMs’ (median) predictions.
@@
Dataset Model Brier Postive Brier ECE ACE
Uncal. Calib. Uncal. Calib. Uncal. Calib. Uncal. Calib.
MATH500 Llama-3.2-1B 0.2414 0.0692 0.2226 0.0472 0.2791 0.0942 0.2791 0.1840
Llama-3.1-8B 0.2045 0.1210 0.1771 0.0994 0.2368 0.1515 0.2368 0.1876
Qwen-2.5-1.5B 0.1541 0.1271 0.1305 0.1072 0.1554 0.1414 0.1554 0.1462
Qwen-2.5-7B 0.1008 0.0818 0.0846 0.0527 0.0981 0.0864 0.0981 0.1227
R1-Llama-8B 0.1614 0.0888 0.1140 0.0546 0.1311 0.0905 0.1311 0.1094
R1-Qwen-7B 0.1480 0.0828 0.1056 0.0578 0.1095 0.0857 0.1095 0.1130
AIME24-25 Llama-3.2-1B 0.1936 0.0029 0.1918 0.0005 0.2364 0.0041 0.4921 0.1306
Llama-3.1-8B 0.2274 0.0414 0.2227 0.0354 0.2839 0.0782 0.4580 0.3988
Qwen-2.5-1.5B 0.3302 0.0727 0.3220 0.0528 0.4007 0.0865 0.3936 0.2889
Qwen-2.5-7B 0.2894 0.0721 0.2820 0.0657 0.3547 0.0982 0.3892 0.2829
R1-Llama-8B 0.3846 0.0782 0.3712 0.0296 0.4764 0.0761 0.3614 0.1566
R1-Qwen-7B 0.4144 0.0694 0.4018 0.0338 0.5078 0.0689 0.3680 0.1261
Significance of calibration. Table 1 presents numerical evidence supporting our earlier claim that PRMs frequently suffer from poor calibration and tend to overestimate. Nevertheless, our proposed calibration approach effectively mitigates these calibration issues. Notably, the approach substantially improves results on out-of-distribution datasets, underscoring its effectiveness across diverse conditions. This finding further indicates that calibrated PRMs are better at differentiating more challenging questions or reasoning tasks from those with higher certainty.
@@ 5.2 Calibrated Reward Enables Instance-Adaptive Scaling
We demonstrate IAS by applying it to two simple yet widely used inference scaling techniques—best-of-$N$ and $N$-beam search on steps.
Instance-adaptive BoN. We evaluate the BoN framework ($N=64$) using both the uncalibrated and calibrated PRMs, comparing their performance against our proposed IAS strategy (BoN+IAS). We adopt a conservative setting with $C=0.99$ and $\beta=0.1$ (see Appendix E.4 for an ablation study). For a fair comparison, we use the calibrated PRM to determine $N$ and the original PRM for ranking after adaptive sampling. We report the accuracy of each method and the relative computational cost associated with IAS, the average number of samples per question normalized by $N$.
@@
Table 2: Comparison between the best-of-$N$ method using a fixed-$N$ strategy (BoN) and our proposed instance-adaptive sampling strategy (BoN+IAS). We report both accuracy and normalized computational cost, measured by the average number of samples per question normalized by $N=64$: $\mathrm{Budget}=N_{\mathrm{IAS}}/N$. Relative improvements over Pass@1 accuracy are highlighted in light blue. Our IAS with calibrated PRMs achieves substantial compute savings without significant performance loss. Crucially, the effectiveness of IAS depends strongly on PRM calibration quality: uncalibrated PRMs tend to overestimate success probabilities, causing overly optimistic downsampling. Moreover, uncalibrated PRM’s model-independent design prevents it from offering adaptive strategies for different LLMs, resulting in further degradation for weaker Llama models.
@@
Dataset Model Pass@1 BoN BoN+IAS Budget BoN+IAS Budget
@@
Table 3: Comparison between the beam search (BS) method using a fixed-$N$/$M$, and our proposed adaptive sampling strategy (IASoK and IASoM) using calibrated PRMs. We report both accuracy and computational cost (budget), measured by the average number of LLM generations per question normalized by that of a fixed-budget BS baseline. Relative improvements over the Pass@1 accuracy are highlighted in light blue. The proposed IAS strategy yields substantial computational savings without significant performance loss.
@@
Dataset Model Pass@1 BS BS+IASoK Budget BS+IASoM Budget
@@
Figure 4: We illustrate average accuracy across test points of varying difficulty levels (1: easy to 5: hard). Results from the fixed-$N$ baseline and our instance-adaptive sampling (IAS) method are shown as dashed lines and stars, respectively. As shown, IAS dynamically adjusts sampling based on problem difficulty in MATH500, allocating more samples to harder tasks.
@@ 6 Conclusion and Limitations
This paper introduces a calibration strategy that enables existing PRMs to produce more reliable success probability estimates. We present a practical use case where well-calibrated PRMs facilitate instance-adaptive inference-time scaling. An interesting path to investigate further is to explore fine-tuning strategies for better PRM calibration that enable broader generalization across datasets and quantile levels. Our empirical study centers on mathematical reasoning tasks, but it would be valuable to extend it to other applications, such as code generation and LLM agents. It would also be interesting to explore IAS in regimes other than BoN/BS. We hope this work establishes a new frontier for reliable reward models and opens a path to more cost-effective, adaptive scaling across domains.

### v2

@@ 4 Instance-Adaptive Inference-Time Scaling
Existing inference-time scaling methods typically allocate a fixed sampling budget NN—for example, by generating $N$ complete reasoning trajectories in BoN. However, this approach can be inefficient: for challenging tasks, $N$ samples may be insufficient to produce a correct answer, while for easier ones, it may waste computation. Just as humans allocate more effort to harder problems and less to simpler ones, it is desirable for LLMs to adjust their compute usage adaptively. Note that in the present work, we (perhaps erring on the side of optimism) assign the maximum budget to the hardest problems, but in principle, we could also choose to say “I don’t know” or route to a more capable model if the probability of success is too low (difficulty too high).
@@
Fix a question–prefix pair and let $p\in[0,1]$ be the success probability (see Definition 1), conditioned on the current generation, that a single continued trajectory sampled from the language model is correct. For a target probability $C\in(0,1)$ we define the independent‑sampling sample‑complexity
[…]
@@
In other words, if $\mathsf{PRM}$ could perfectly distinguish correct from incorrect reasoning, then selecting the best trajectory among $N_{\mathrm{IAS}}(p,C)$ samples (i.e., “best-of-$N_{\mathrm{IAS}}$”) guarantees an average accuracy of (at least) $C$ for questions whose per-trajectory success probability is $p$.
@@
As noted previously, PRMs frequently overestimate LLMs’ success probabilities—particularly for weaker models or on hard, out-of-distribution queries (Figure 2). Relying on these inflated scores leads to suboptimal inference-time scaling by selecting an excessively small $N_{\mathrm{IAS}}$, whereas calibration mitigates these pitfalls and enables principled, cost-effective decision making.
@@ 4.1 IAS: Calibrated Reward-based Instance-Adaptive Scaling
BoN+IAS. Rather than drawing a fixed sample size $N$ for best-of-$N$ decoding, our IAS framework adaptively determines the number of trajectories to generate based on the estimated difficulty of each problem. Specifically, we compute the minimum number of samples $N_{\mathrm{IAS}}$ needed to achieve the target correctness level $C$ given the PRM’s estimated success probability $\hat{r}^{(\beta)}$:
$N_{\mathrm{IAS}}\triangleq\min\{\lceil N_{\mathrm{IAS}}(\hat{r}^{(\beta)},C)\rceil\;,\;N_{\max}\},$
where $N_{\max}$ is the maximum budget constraint. For problems with high estimated success probability, $N_{\mathrm{IAS}}$ will be small, while challenging problems receive larger budgets up to $N_{\max}$.
BS+IAS-of-$M$ (with fixed $K$). In beam search with a fixed beam width $K$, we can adaptively determine how many continuations to sample per prefix at each step. Suppose at a given beam-search step we have $K$ candidate prefixes. Let $\hat{r}^{(\beta)}{\mathrm{min}}$ denote the minimum calibrated success probability among these $K$ prefixes. To ensure at least one correct completion across all $K$ prefixes with confidence level $C$, IAS computes the number of samples per prefix as:
$M_{\mathrm{IAS}}\triangleq\min\Bigl\{\Bigl\lceil\tfrac{N_{\mathrm{IAS}}\!\bigl(\hat{r}^{(\beta)}_{\mathrm{min}},C\bigr)}{\,K\,}\Bigr\rceil\;,\;M_{\max}\Bigr\}.$
where $M_{\max}$ is the maximum number of expansions allowed per prefix. By using the most pessimistic estimate $\hat{r}^{(\beta)}_{\mathrm{min}}$, we ensure sufficient sampling even for the most challenging prefix in the beam.
BS+IAS-of-$K$ (with fixed $M$). Conversely, when the number of expansions per prefix $M$ is fixed, we can adaptively determine the beam width itself. After generating $M$ continuations for each of $K$ candidate prefixes and ranking them by their calibrated rewards $\hat{r}^{(\beta)}$, IAS determines how many top-ranked prefixes to retain:
$K_{\mathrm{IAS}}\triangleq\min\left(\{K_{\max}\}\cup\{k\mid N_{\mathrm{IAS}}(\hat{r}_{k}^{(\beta)},C)\leq k\times M,\quad 1\leq k\leq K_{\max}\}\right)$
where $\hat{r}_{k}^{(\beta)}$ is the calibrated reward of the $k$-th best prefix and $K_{\max}$ is the maximum allowed beam width. This ensures that the total budget $K_{\mathrm{IAS}}\times M\;\geq\;N_{\mathrm{IAS}}\!\bigl(\hat{r}_{K_{\mathrm{IAS}}}^{(\beta)},C\bigr)$, maintaining the target confidence level $C$ that at least one prefix will lead to a correct answer.
@@
[…]
where $\tilde{\beta}\geq\beta$ depends on $\mathcal{V}_{n}$ (see Appendix B.2.1 for its specification).
@@ 5.1 Fine-Tuning PRMs for Better Calibration
We evaluate calibration errors of off-the-shelf PRMs and then show how our fine-tuning strategy reduces these errors. To start with, we construct a calibration dataset by randomly sampling 500 questions from the MATH training set. We assess calibration performance using standard metrics: Brier score [3], positive Brier score (i.e., the mean square of overestimation error, $\max\{\hat{y}-y,0\}$), AdaptiveCE [34], ECE [32], and AverageCE [33] between ground-truth Monte Carlo success rates and the PRMs’ (median) predictions.
Dataset Model Brier PosBrier AdaptiveCE ECE AverageCE
Uncal. Calib. Uncal. Calib. Uncal. Calib. Uncal. Calib. Uncal. Calib.
MATH500 Llama-3.2-1B 0.2414 0.0692 0.2226 0.0472 0.2830 0.0811 0.2791 0.0942 0.3130 0.1840
Llama-3.1-8B 0.2045 0.1210 0.1771 0.0994 0.2625 0.1674 0.2368 0.1515 0.2048 0.1876
Qwen-2.5-1.5B 0.1541 0.1271 0.1305 0.1072 0.2176 0.1545 0.1554 0.1414 0.1229 0.1462
Qwen-2.5-7B 0.1008 0.0818 0.0846 0.0527 0.1459 0.0999 0.0981 0.0864 0.0920 0.1227
R1-Llama-8B 0.1614 0.0888 0.1140 0.0546 0.2505 0.1128 0.1311 0.0905 0.1517 0.1094
R1-Qwen-7B 0.1480 0.0828 0.1056 0.0578 0.2413 0.1066 0.1095 0.0857 0.1957 0.1130
AIME24-25 Llama-3.2-1B 0.1936 0.0029 0.1918 0.0005 0.2364 0.0108 0.2364 0.0041 0.4921 0.1306
Llama-3.1-8B 0.2274 0.0414 0.2227 0.0354 0.2839 0.0862 0.2839 0.0782 0.4580 0.3988
Qwen-2.5-1.5B 0.3302 0.0727 0.3220 0.0528 0.4007 0.1054 0.4007 0.0865 0.3936 0.2889
Qwen-2.5-7B 0.2894 0.0721 0.2820 0.0657 0.3547 0.0982 0.3547 0.0982 0.3892 0.2829
R1-Llama-8B 0.3846 0.0782 0.3712 0.0296 0.5259 0.1275 0.4764 0.0761 0.3614 0.1566
R1-Qwen-7B 0.4144 0.0694 0.4018 0.0338 0.5575 0.0898 0.5078 0.0689 0.3680 0.1261
@@
Significance of calibration. Table 1 presents numerical evidence supporting our earlier claim that PRMs frequently suffer from poor calibration and tend to overestimate. In contrast, our proposed calibration approach effectively mitigates these calibration issues. Additionally, the approach substantially improves results on out-of-distribution datasets, underscoring its effectiveness across diverse conditions. This finding further indicates that calibrated PRMs are better at differentiating more challenging questions or reasoning tasks from those with higher certainty.
@@ 5.2 Calibrated Reward Enables Instance-Adaptive Scaling
We demonstrate IAS by applying it to two simple yet widely used inference scaling techniques: best-of-$N$ and $N$-beam search on steps.
Instance-adaptive BoN. We evaluate the BoN framework ($N_{\max}=64$) using both the uncalibrated and calibrated PRMs, comparing their performance against our proposed IAS strategy (BoN+IAS). We adopt a conservative setting with $C=0.99$ and $\beta=0.1$ (see Appendix E.5 for an ablation study). For a fair comparison, we use the calibrated PRM to determine $N_{\text{IAS}}$ and the original PRM for ranking after adaptive sampling. We report the accuracy of each method and the relative computational cost associated with IAS, the average number of samples per question normalized by $N_{\max}$.
@@
Table 2: Comparison between the best-of-$N$ method using a fixed-NN strategy (BoN) and our proposed instance-adaptive sampling strategy (BoN+IAS). We report both accuracy and relative computational cost (budget), measured by the average number of samples per question normalized by $N=64$: $\mathrm{Budget}=N_{\mathrm{IAS}}/N$. Relative improvements over Pass@1 accuracy are highlighted in light blue. Our IAS with calibrated PRMs achieves substantial compute savings without significant performance loss. Crucially, the effectiveness of IAS depends strongly on PRM calibration quality: uncalibrated PRMs tend to overestimate success probabilities, causing overly optimistic downsampling. Moreover, uncalibrated PRM’s model-independent design prevents it from offering adaptive strategies for different LLMs, resulting in further degradation for weaker Llama models.
@@
Dataset Model Pass@1 BoN BoN+IAS Budget Ratio BoN+IAS Budget Ratio
@@
Table 3: Comparison between the beam search (BS) method using a fixed-NN/$M$, and our proposed adaptive sampling strategy (IASoK and IASoM) using calibrated PRMs. We report both accuracy and relative computational cost (budget) to the baseline, measured by the average number of LLM generations per question normalized by that of a fixed-budget BS baseline. Relative improvements over the Pass@1 accuracy are highlighted in light blue. The proposed IAS strategy yields substantial computational savings without significant performance loss.
@@
Dataset Model Pass@1 BS BS+IASoK Budget Ratio BS+IASoM Budget Ratio
@@
Figure 4: We illustrate average accuracy across test points of varying difficulty levels (1: easy to 5: hard). Results from the fixed-NN baseline and our instance-adaptive sampling (IAS) method are shown as dashed lines and stars, respectively. As shown, IAS dynamically adjusts sampling based on problem difficulty in MATH500, allocating more samples to harder tasks.
@@ 6 Conclusion and Limitations
This paper introduces a calibration strategy that enables PRMs to produce more reliable estimates of success probability: the likelihood that a partial reasoning step (or initial question) will lead to a correct final answer when completed by a given policy LLM. Through quantile regression, our approach provides conservative lower bounds rather than overoptimistic point estimates, addressing a key limitation of standard PRMs.
We demonstrate a practical application of well-calibrated PRMs in instance-adaptive inference-time scaling (IAS), where computational budget is allocated based on estimated success probability. By providing reliable estimates, our calibrated PRMs enable more efficient resource allocation, investing more computation on challenging problems while avoiding waste on easier ones.
Future work could explore fine-tuning strategies for better generalization across datasets and quantile levels, extend this approach beyond mathematical reasoning to domains like code generation and LLM agents, and investigate IAS in settings beyond best-of-N and beam search. We hope this work establishes a foundation for more reliable reward models and cost-effective adaptive scaling.